if points A,B,C on a coordinate line have cooridnates -5,3,and 10 respectively, find d(A,C) + d(B,C).
Answer by venugopalramana(3286)

if points A,B,C on a coordinate line have cooridnates -5,3,and 10 respectively, find d(A,C) + d(B,C).
THEY ARE ALL ON A COORDINATE LINE..THAT IS AXIS...SO ....
AC=10-(-5)=15
BC=10-3=7
AC+BC=15+7=22